The nucleus of an atom consists of protons and neutrons (no electrons). A nucleus of a carbon-12 isotope contains six protons and six neutrons, while nitrogen-14 nucleus comprises seven protons and seven neutrons. You perform a nuclear physics experiment in which you bombard nitrogen-14 nuclei with very high speed carbon-12 nuclei emerging from a particle accelerator. The nucleus of an atom consists of protons and neu
Answer #1

here

let the number of electrons is n

as the charge is conserved during the collision

charge of antiproton = -1

charge on position = + 1

charge on proton = + 1

charge on electron = - 1

total initial charge = 6 + 7 = 13

initial charge = final charge

let the number of electrons be n

16* 1 + 3 * ( -1) + 7 * ( 1) + n * ( -1 ) = 13

n = 7

the number of emitted electrons is 7 electrons